How much 2% acid must be added to 12% acid to get 50 ml of 10% acid?

Answer by richwmiller(17219)

a+b=50,
0.02*a+0.12*b=0.1*50
a=50-b
0.02*(50-b)+0.12*b=5
1-0.02b+0.12*b=5
0.1*b=4
b=40
a=50-b
a=10 b=40
check
0.02*10+0.12*40=0.1*50
0.2+4.8=5
5=5
ok